# Mutation-Guided Unit Test Generation with a Large Language Model

### v5

@@ III-C Prompting with Mutation Feedback
As illustrated in Figure 6 and the example in Section II-B, the prompt constructed by MutGen includes a natural language code summarization (descdesc), the source code excluding comments (srcsrc), and the associated mutation feedback ($R$). In Figure 4, these correspond to the values of the prompt variables: descdesc is represented as {{summary}}, srcsrc as {{source code without comments}}, and $\mathcal{R}$ as the content under ## Mutation Feedback. In addition, MutGen explicitly specifies the expected output format from the LLM, as detailed in the prompt example Figure 4 shown in Section II-B.
III-D Fixing Execution Errors
After the initial generation stage, some test cases may fail due to errors in their code. MutGen further introduces a fixing step that applies a targeted prompt, as illustrated in the example prompt shown in Figure 5, to repair them.
We categorize these failures into two types:
  • Assertion Failures: Test cases that can pass compilation but fail to execute due to violated assertions or incorrect outputs.
  • Compilation Errors: Test cases that cannot be compiled due to syntax errors, missing types, unresolved references, or similar issues.
As shown in prior work [yuan2024evaluating], advanced LLMs are generally capable of producing syntactically correct test cases. However, they often suffer from runtime errors, among which assertion failures are widespread. In this paper, which targets unit testing scenarios, we focus on addressing assertion failures. To this end, MutGen incorporates four fix strategies into the prompt: Direct Invocation, Handling AssertEquals Mismatches, Fixing assertTrue/assertFalse Errors, and Handling Ambiguous Cases. For potential compilation errors, we employ a fifth strategy, which proactively avoids these issues during the process described in Section III-E. For other errors, MutGen adopts a sixth strategy, instructing the LLM to actively repair them rather than ignoring them. All fix strategies are listed in the prompt shown in Figure 5.
We partition the set of failures into compilation and assertion failures such that
$\mathcal{F}=\mathcal{F}_{\text{comp}}\cup\mathcal{F}_{\text{assert}},\quad\mathcal{F}_{\text{comp}}\cap\mathcal{F}_{\text{assert}}=\emptyset.$
Each failing test case is associated with its corresponding error message, forming a tuple $f_{i}=(t_{i},e_{i})$, where $t_{i}$ is the original test case and $e_{i}$ is the compiler or runtime error message.
We define a prompt construction function:
@@ III-D Fixing Execution Errors
which creates a structured natural language prompt, as illustrated in Figure 5.
We define the specialized fix functions as:
@@
$\mathcal{T}^{\prime\prime}$ $=\{\text{Fix}_{\text{assert}}(f_{i})\mid f_{i}\in\mathcal{F}_{\text{assert}}\}.$
We denote the final set of fixed test cases as:
@@
III-E Pushing the Limit: Iterative Generation
@@
The process continues until convergence (i.e., no significant improvement in mutation score) or until a predefined iteration limit is reached. Once convergence is reached, the remaining live and uncovered mutants provide insights into mutation operators and program constructs that remain challenging for the LLM, which we analyze in Section IV-B.
IV Evaluation
As previously discussed, MutGen is designed to enhance the fault-detection capability of LLMs in automated test generation by incorporating mutation feedback. To evaluate its effectiveness, we compare MutGen with three baselines: EvoSuite, the state-of-the-art search-based technique; EvoSuitemut, a variant that replaces coverage-based fitness functions with strong mutation as described in [fraser2015achieving]; and Genvanilla, a prompting-based approach that utilizes an LLM without incorporating mutation feedback and leveraging the comments provided in the program under test through summarization (as illustrated in Figure 4). We adopt Llama-3.3 70B as the base model in our evaluation due to its recent release, strong effectiveness on code-related tasks [grattafiori2024llama], and open-source accessibility, which ensures reproducibility and ease of deployment in both research and practical scenarios.
@@ IV Evaluation
To sum up, our evaluation addresses the following research questions.
@@
IV-A Experiment Setup
Subjects. In our study, we primarily focus on method-level subjects, where each subject is a single Java class containing only one target method. We generate unit test cases specifically for this method. We use a widely adopted benchmark dataset that has been used in prior publications [siddiq2024using, zheng2023codegeex, tian2025fixing, islam2405mapcoder], and we also create an additional dataset to enable a more diverse evaluation. More specifically, we include the following subjects:
@@ IV-A Experiment Setup
Following prior work [islam2405mapcoder, siddiq2024using, tian2025fixing, zheng2023codegeex], we adopt the HumanEval dataset [chen2021evaluating, athiwaratkun2022multi] as one of our evaluation benchmarks. Specifically, we utilize all 160 Java subjects from the dataset. For each subject, we apply EvoSuite with its default configuration to automatically generate a test suite. We then evaluate the mutation score of the generated test suite using PITest [pitest]. Subjects where MutGen and EvoSuite both achieved a 100% mutation score are discarded, resulting in 104 Java subjects for our evaluation. Indeed, these subjects do not help us distinguish the two approaches.
@@
We search GitHub using the keyword "Leetcode Java" and select the top two repositories based on the Best match sorting criterion. These repositories contain Java solutions to algorithmic problems from LeetCode [leetcode], which are categorized into three difficulty levels: Easy, Medium, and Hard. We exclude all solutions corresponding to Easy problems. We collect all Java source files for the remaining Medium and Hard problems and organize them into a unified package structure following the convention used in HumanEval-Java. We use the Maven build system to compile the project and remove any subjects that failed to compile. For each valid subject, we apply EvoSuite with its default settings to generate a test suite. Subjects for which the generated test suite achieved a 100% mutation score by MutGen and EvoSuite are once again discarded. Finally, for computational feasibility, we randomly select 50 subjects from each of the Medium and Hard sets, resulting in a total of 100 Leetcode-Java subjects used in our evaluation.
The average number of lines of code in HumanEval-Java methods is 41, with an average cyclomatic complexity (CC) of 4.90, which is close to the average CC of 5.46 reported for typical Java methods [cc]. This indicates that HumanEval-Java exhibits representative control-flow complexity for standard Java methods. In contrast, methods in Leetcode-Java are generally more complex, with an average CC of 7.88. These statistics demonstrate that Leetcode-Java contains methods with complexity levels comparable to or exceeding those found in real-world Java code [ccpaper], making it both representative and more challenging than HumanEval-Java.
@@
A mutant is considered killed if at least one generated test case causes a behavioral difference between the original program and its mutated version, typically resulting in a failed assertion or exception. In addition to the mutation score, we also report branch coverage and line coverage, in line with previous work [alagarsamy2024a3test, chen2024chatunitest].
@@
Process. To address RQ1, we use all subjects from both datasets. We begin by generating test cases with both Genvanilla and MutGen through a single end-to-end execution, recording the execution time and the number of tokens consumed. We then apply EvoSuite and EvoSuitemut to generate test cases for each subject, setting their time budget to the larger average processing times of MutGen on the two datasets. It is important to note that, unlike prior work [foster2025mutation], our approach generates tests entirely from scratch, without relying on any existing test cases. We evaluate each test suite using two structural coverage metrics, branch coverage and line coverage, measured by the JaCoCo tool [jacoco]. Additionally, we use PITest [pitest] with its DEFAULTS mutation operator group to measure the mutation score of each test suite. The DEFAULTS group provides a set of operators commonly used in the literature [fraser2015achieving, wang2024comprehensive] and, according to the PITest official documentation [pitest], are not easy to detect and minimize the number of equivalent mutants. While PITest also offers a STRONGER operator group, it differs from DEFAULTS by only two additional mutation operators, whose applicability is rather limited. For example, the Remove Conditionals operator only targets equality checks. As for the ALL group provided by PITest, the extra mutation operators in this group generally generate mutants that are relatively easy to kill, making them less informative for our evaluation. Note that although PITest is designed for Java, its mutation operators can be applied to other languages as well [pitest, wang2024comprehensive, jia2010analysis]. To assess the statistical significance of the improvements achieved by our approach in terms of effectiveness, we performed non-parametric hypothesis testing using Vargha and Delaney’s A12 statistic [arcuri2011practical, vargha2000critique]. The A12 effect size was computed based on the mutation score, which ranges from 0 to 1: a value of 0.5 indicates no difference (i.e., results are due to chance), while a value greater than 0.5 suggests that MutGen is more likely to outperform the compared approach. Conversely, a value less than 0.5 indicates the opposite. We also conducted a paired-sample Wilcoxon signed-rank test on both coverage and mutation score to evaluate whether MutGen achieves statistically significant improvements on these metrics.
@@
To address RQ2, we analyze the mutation operators involved in the evaluation. As a preliminary experiment, we first randomly select 10 subjects from each dataset and perform seven rounds of end-to-end test generation. In each round, MutGen generates additional test cases from previously generated ones. By tracking mutation scores across iterations, we observe that they tend to converge after the fourth iteration. Therefore, we set the maximum number of repetitions to four in our experiments. We collect results at the fourth iteration, and analyze the distribution of mutation operators across all generated mutants and compute the killing ratio for each operator, that is, the proportion of killed mutants relative to the total number of mutants produced by that operator. This allows us to identify which mutation operators tend to produce mutants that are difficult or easy for LLM-generated test cases to kill. In our evaluation, mutation testing is conducted using PITest, which employs 11 mutation operators as described in [mutops]. For mutation operators with low killing ratios, we investigate why LLMs struggle to kill these mutants. Due to space constraints, we do not present a detailed analysis for each mutation operator or for all live and uncovered mutants. Instead, we analyze a subset of cases and identify several factors that contribute to LLMs’ inability to kill certain mutants. Nevertheless, our findings offer valuable insights and highlight important directions for future research in this area.
@@
To address RQ3, we conduct an ablation study to evaluate the contribution of each core component in MutGen. Specifically, MutGen consists of three main components: (1) extracting and incorporating code summarization into prompts, (2) applying the fixing step to repair invalid test cases, and (3) extracting and incorporating mutation feedback into prompts. In this study, we disable one component at a time, resulting in three variants MutGen-S (using the original code comments instead of summarization), MutGen-F (omitting the fixing step), and MutGen-MF (excluding mutation feedback). For each variant, we perform four iterations of test case generation and measure mutation score to assess the impact of each component on the effectiveness of MutGen.
@@
Implementation. We describe important implementation details of MutGen and our evaluation setup. All subjects used in the evaluation are built using the Maven build system. We follow official guidelines to integrate EvoSuite, JaCoCo, and PITest into the Maven pipeline for automated test generation, code coverage measurement, and mutation testing, respectively. Our approach, MutGen, is implemented in Python and built on top of the Mutahunter framework [mutahunter]. Specifically, we reuse components from Mutahunter for prompt handling, LLM interfacing, and response processing. We use Ollama [ollama] to locally deploy the Llama-3.3 70B model [grattafiori2024llama] with the default configuration, setting the temperature to 0.0 for more deterministic outputs. As our main baseline, EvoSuite officially and stably supports JUnit 4, while its support for JUnit 5 is currently limited and experimental. Therefore, JUnit 4 remains the recommended choice for reliable test generation. To ensure a fair comparison, we adopt JUnit 4 in all experimental settings.
@@ IV-B Results and Analysis
On HumanEval-Java, for both line coverage and branch coverage, MutGen outperforms EvoSuite, EvoSuitemut and Genvanilla. On Leetcode-Java, we see the opposite: EvoSuite achieves the highest coverage scores, surpassing both LLM-based techniques. For both datasets, however, differences are relatively small, below three percentage points. Although MutGen is primarily designed to improve mutation score, the results also demonstrate that it consistently outperforms Genvanilla in terms of line and branch coverage. This supports previous observations  [foster2025mutation] that improvements in mutation score can also lead to enhanced code coverage.
@@
Regarding the mutation score, PITest generates an average of 11 and 19 mutants per subject on HumanEval-Java and Leetcode-Java, respectively, resulting in a total of 1,144 and 1,900 mutants across the two datasets. MutGen achieves mutation scores of 89.5% and 89.1% on the two datasets. Compared to EvoSuite, this represents large improvements of 28.8% and 51.3%, respectively, corresponding to around 20 and 30 percentage points. Compared to Genvanilla, MutGen also significantly improves the mutation score, though to a lesser extent. Additionally, MutGen also surpasses EvoSuitemut, which specifically aims to maximize the mutation score, to an even larger extent. Indeed, surprisingly, we observe that EvoSuite slightly outperforms EvoSuitemut. This might be due to the latter being developed 10 years ago [fraser2014large] and receiving no further maintenance. For example, we found that the mutation operators used by EvoSuitemut are limited, resulting in poor performance of the generated test cases when evaluated by PITest. Another tentative explanation is that, to ensure fair comparisons, recall that we set the time budget to the maximum used by MutGen, but a larger budget may be necessary for EvoSuitemut.
@@
Furthermore, we report the A12 effect size [arcuri2011practical] on mutation scores in the third-to-last column of Table I. For the comparison between EvoSuite and MutGen, the A12 values are 0.759 and 0.899, indicating that MutGen has a 75.9% and 89.9% probability of achieving higher mutation scores than EvoSuite on HumanEval-Java and Leetcode-Java, respectively. Similarly, when compared with Genvanilla, the A12 values are 0.650 and 0.734, suggesting that MutGen outperforms Genvanilla with 65.0% and 73.4% probabilities on the same datasets. Additionally, based on the results of the paired-sample Wilcoxon signed-rank test, we found that MutGen achieves statistically significant improvements ($\alpha$ = 0.05) in mutation scores compared with EvoSuite, EvoSuitemut, and Genvanilla. However, there is no statistically significant difference in coverage, as most subjects already achieve near 100% coverage across all approaches.
@@
In terms of overhead, MutGen consumes fewer tokens on HumanEval-Java but more tokens on Leetcode-Java. The Leetcode-Java subjects are solutions to medium- and hard-level algorithmic problems, with more complex logic than those in HumanEval-Java. As a result, PITest generates more mutants for them (19 vs. 11), and Llama-3.3 requires more test cases to kill all mutants. Compared with Genvanilla, EvoSuite, and EvoSuitemut, in Leetcode-Java, MutGen consumes more tokens while achieving higher mutation scores as well as better line and branch coverage. In contrast, on HumanEval-Java, MutGen consumes fewer tokens than EvoSuite and EvoSuitemut while still achieving superior performance in both coverage and mutation score. For the baseline approaches (i.e., EvoSuite and EvoSuitemut), a possible explanation for their higher token consumption on HumanEval-Java is that these programs contain more lines of code on average, whereas LeetCode-Java programs exhibit higher cyclomatic complexity. Longer programs tend to result in larger or more verbose test suites generated by these search-based techniques, leading to increased token usage at the test level. Interestingly, we observe an opposite trend for our LLM-based approach. On LeetCode-Java, MutGen consumes more tokens than on HumanEval-Java, despite the latter having longer programs on average. We hypothesize that this behavior is related to the use of mutation feedback in the prompt. Specifically, LeetCode-Java has a larger number of mutants per subject, leading to more uncovered or live mutants included in the mutation feedback provided to the LLM. As a consequence, the prompt becomes richer and more informative, encouraging the LLM to generate more extensive test cases to kill these mutants, which, in turn, leads to higher token consumption. However, differences of that scale in the number of tokens do not have practical implications.
@@
Overall, results suggest that MutGen significantly improves mutation scores over alternatives, with large differences, at the expense of slightly lower code coverage compared to EvoSuite for the most complex subject (Leetcode-Java). This is expected, as EvoSuite is designed to maximize coverage. For the other subject (HumanEval-Java), the reverse trend is observed. Even when EvoSuite is specifically configured as EvoSuitemut to maximize mutation scores [fraser2014large], it performs slightly worse than EvoSuite on both datasets. Therefore, to optimize mutation scores and, thus, fault detection, we recommend using MutGen, though the impact on coverage is unclear and likely negligible in practice.

### v6

@@ II-D Prompting with Mutation Feedback
As illustrated in Figure 2 and the example in Section II-A, the prompt constructed by MutGen includes three components: a natural language code summarization (descdesc), the source code excluding comments (srcsrc), and the associated mutation feedback ($\mathcal{R}$). In Figure 5, the former two correspond to the prompt variables: descdesc is represented as {{summary}}, and srcsrc as {{source code without comments}}.
Regarding the content under ## Mutation Feedback, we first provide a sentence-level instruction to the LLM. The LLM is provided with live mutants and is prompted to generate tests that are more diverse for each mutant. Following this sentence-level prompt, the set of mutation feedback entries, $\mathcal{R}$, is provided as defined in Section II-C.
In addition, MutGen explicitly specifies the expected output format from the LLM, as illustrated in the prompt example shown in Figure 5.
Figure 5: Prompt used for generation
II-E Fixing Execution Errors
After the initial generation stage, some test cases may fail due to execution errors. MutGen introduces a fixing step that leverages a structured prompt, as illustrated in Figure 6, to guide the LLM in repairing these failures. In the example shown, two generated test cases initially fail to execute due to runtime errors. By applying the fixing step, MutGen successfully repairs one of the failing test cases, resulting in a 13% increase in mutation score.
We categorize execution failures into two types:
  • Assertion Failures: Test cases that compile successfully but fail at runtime due to violated assertions or incorrect outputs.
  • Compilation Errors: Test cases that fail to compile due to syntax errors, missing types, unresolved references, or similar issues.
To address these failures, MutGen constructs a structured prompt for each failing test case:
@@ II-E Fixing Execution Errors
where $t_{i}$ is the original test case and $e_{i}$ is the associated compiler or runtime error message. The prompt explicitly instructs the LLM:
”You are processing execution failures. Please match only one of these guidelines and then try to correct the failed tests according to the given error messages.”
The prompt then enumerates six targeted fix strategies:
  1. 1.
    Direct invocation errors: to address incorrect invocation of assertion functions.
  2. 2.
    assertEquals mismatches: to address oracle value inconsistencies in assertEquals assertions.
  3. 3.
    assertTrue/assertFalse errors: to address binary oracle inconsistencies in assertTrue or assertFalse assertions.
  4. 4.
    Ambiguous assertion references: to address ambiguity in assertEquals calls; explicit type casts are applied to clarify the parameter types.
  5. 5.
    Method name conflicts: to address duplicate method names within the test class.
  6. 6.
    Other runtime or compilation errors: to address errors not covered by the above strategies.
Detailed instructions for these strategies, along with a fixing example, are provided in Figure 6.
We define specialized fix functions for each failure type:
@@
$\mathcal{T}^{\prime\prime}$ $=\{\text{Fix}_{\text{assert}}(f_{i})\mid f_{i}\in\mathcal{F}_{\text{assert}}\},$
and the final set of fixed test cases is:
@@
This prompt-based fixing strategy ensures that the LLM systematically addresses different types of execution failures, improving test execution success and contributing to higher mutation scores.
Prompt Used for Fixing ## Failed Test The following test failed in the test suite: {{ failed_test with error message }} You are processing execution failures. Please match only one of these guidelines and then try to correct the failed tests according to the given error messages. 1. Direct Invocation: • Always use assertion functions directly (e.g., assertEquals(value1, value2)) instead of prefixed calls like Assert.assertEquals(value1, value2). 2. Handling assertEquals Mismatches: • If an error follows the pattern: • test_file_name:line_number: expected:error_value but was:correct_value • And involves assertEquals, replace the expected value with the correct one. • Example: assertEquals(error_value, xxx); → assertEquals(correct_value, xxx); 3. Fixing assertTrue / assertFalse Errors: • If an error follows the pattern: • classname:line_number:null • And involves assertTrue or assertFalse, swap the function name while keeping all parameters unchanged. • Example: assertTrue(xxx); → assertFalse(xxx); 4. Handling ambiguous • If an error has the pattern: • reference to assertEquals is ambiguous • And involves assertEquals, add type casts to both parameters based on their respective types. • Example: assertEquals(expectedValue, actualValue); → assertEquals((long) expectedValue, (long) actualValue); 5. Method name conflicts • If an error follows the pattern: method is already defined in the class, • only append a number to the method name. • Example: public void test_null() → public void test_null1(); 6. Other errors • Try to fix the errors according to the error messages. ## Output Format in YAML The same applied as in Figure 5.
Figure 6: Prompt used for fixing
II-F Pushing the Limit: Iterative Generation
@@
The process continues until convergence (i.e., no significant improvement in mutation score) or until a predefined iteration limit is reached. Once convergence is reached, the remaining live and uncovered mutants provide insights into mutation operators and program constructs that remain challenging for the LLM, which we analyze in Section III-B.
To illustrate the effectiveness of this iterative process, we revisit the running example introduced earlier. After two iterations, for the running example, MutGen achieves a 100% mutation score. In contrast, as demonstrated in Section III, for this subject in HumanEval-Java, using the vanilla prompt introduced at the beginning of this section, the LLM produces a test suite with a 53% mutation score, which remains unchanged after four iterations.
III Evaluation
As previously discussed, MutGen is designed to enhance the fault-detection capability of LLMs in automated test generation by incorporating mutation feedback. To evaluate its effectiveness, we compare MutGen with three baselines: EvoSuite, the state-of-the-art search-based technique; EvoSuitemut, a variant that replaces coverage-based fitness functions with strong mutation as described in [27]; and Genvanilla, a prompting-based approach that utilizes an LLM without incorporating mutation feedback and leveraging the comments provided in the program under test through summarization (as illustrated in Figure 5). We adopt Llama-3.3 70B as the base model in our evaluation due to its recent release, strong effectiveness on code-related tasks [30], and open-source accessibility, which ensures reproducibility and ease of deployment in both research and practical scenarios.
@@ III Evaluation
In sum, our evaluation addresses the following research questions.
@@
III-A Experiment Setup
Subjects. In our study, we primarily focus on method-level subjects, where each subject is a single Java class containing only one target method. We generate unit test cases specifically for this method. We use a widely adopted benchmark dataset that has been used in prior publications [65, 81, 68, 37], and we also create an additional dataset to enable a more diverse evaluation. More specifically, we include the following subjects:
@@ III-A Experiment Setup
Following prior work [37, 65, 68, 81], we adopt the HumanEval dataset [13, 8] as one of our evaluation benchmarks. Specifically, we utilize all 160 Java subjects from the dataset. For each subject, we apply EvoSuite with its default configuration to automatically generate a test suite. We then evaluate the mutation score of the generated test suite using PITest [60]. Subjects where MutGen and EvoSuite both achieved a 100% mutation score are discarded, resulting in 104 Java subjects for our evaluation. Indeed, these subjects do not help us distinguish the two approaches.
@@
We search GitHub using the keyword "Leetcode Java" and select the top two repositories based on the Best match sorting criterion. These repositories contain Java solutions to algorithmic problems from LeetCode [43], which are categorized into three difficulty levels: Easy, Medium, and Hard. We exclude all solutions corresponding to Easy problems. We collect all Java source files for the remaining Medium and Hard problems and organize them into a unified package structure following the convention used in HumanEval-Java. We use the Maven build system to compile the project and remove any subjects that failed to compile. For each valid subject, we apply EvoSuite with its default settings to generate a test suite. Subjects for which the generated test suite achieved a 100% mutation score by MutGen and EvoSuite are once again discarded. Finally, for computational feasibility, we randomly select 50 subjects from each of the Medium and Hard sets, resulting in a total of 100 Leetcode-Java subjects used in our evaluation.
The average number of lines of code in HumanEval-Java methods is 41, with an average cyclomatic complexity (CC) of 4.90, which is close to the average CC of 5.46 reported for typical Java methods [52]. This indicates that HumanEval-Java exhibits representative control-flow complexity for standard Java methods. In contrast, methods in Leetcode-Java are generally more complex, with an average CC of 7.88. These statistics demonstrate that Leetcode-Java contains methods with complexity levels comparable to or exceeding those found in real-world Java code [66], making it both representative and more challenging than HumanEval-Java.
@@
A mutant is considered killed if at least one generated test case causes a behavioral difference between the original program and its mutated version, typically resulting in a failed assertion or exception. In addition to the mutation score, we also report branch coverage and line coverage, in line with previous work [2, 14].
@@
Process. To address RQ1, we use all subjects from both datasets. We begin by generating test cases with both Genvanilla and MutGen through a single end-to-end execution, recording the execution time and the number of tokens consumed. We then apply EvoSuite and EvoSuitemut to generate test cases for each subject, setting their time budget to the larger average processing times of MutGen on the two datasets. It is important to note that, unlike prior work [24], our approach generates tests entirely from scratch, without relying on any existing test cases. We evaluate each test suite using two structural coverage metrics, branch coverage and line coverage, measured by the JaCoCo tool [38]. Additionally, we use PITest [60] with its DEFAULTS mutation operator group to measure the mutation score of each test suite. The DEFAULTS group provides a set of operators commonly used in the literature [27, 71] and, according to the PITest official documentation [60], are not easy to detect and minimize the number of equivalent mutants. While PITest also offers a STRONGER operator group, it differs from DEFAULTS by only two additional mutation operators, whose applicability is rather limited. For example, the Remove Conditionals operator only targets equality checks. For the ALL group provided by PITest, the additional mutation operators generally generate mutants that are relatively easy to kill, making them less informative for our evaluation. Note that although PITest is designed for Java, its mutation operators can be applied to other languages as well [60, 71, 39]. To assess the statistical significance of the improvements achieved by our approach in terms of effectiveness, we performed non-parametric hypothesis testing using Vargha and Delaney’s A12 statistic [7, 69]. The A12 effect size was computed from the mutation score, which ranges from 0 to 1: a value of 0.5 indicates no difference (i.e., results are due to chance), whereas a value greater than 0.5 indicates that MutGen is more likely to outperform the compared approach. Conversely, a value less than 0.5 indicates the opposite. We also conducted a paired-sample Wilcoxon signed-rank test on both coverage and mutation score to evaluate whether MutGen achieves statistically significant improvements on these metrics.
@@
To address RQ2, we analyze the mutation operators involved in the evaluation. As a preliminary experiment, we first randomly select 10 subjects from each dataset and perform seven rounds of end-to-end test generation. In each round, MutGen generates additional test cases from previously generated ones. By tracking mutation scores across iterations, we observe that they tend to converge after the fourth iteration. Therefore, we set the maximum number of repetitions to four in our experiments. We collect results at the fourth iteration, and analyze the distribution of mutation operators across all generated mutants and compute the killing ratio for each operator, that is, the proportion of killed mutants relative to the total number of mutants produced by that operator. This allows us to identify which mutation operators tend to produce mutants that are difficult or easy for LLM-generated test cases to kill. In our evaluation, mutation testing is conducted using PITest, which employs 11 mutation operators as described in [20]. For mutation operators with low killing ratios, we investigate why LLMs struggle to kill these mutants. Due to space constraints, we do not present a detailed analysis for each mutation operator or for all live and uncovered mutants. Instead, we analyze a subset of cases and identify several factors that contribute to LLMs’ inability to kill certain mutants. Nevertheless, our findings offer valuable insights and highlight important directions for future research in this area.
@@
To address RQ3, we conduct an ablation study to evaluate the contribution of each core component in MutGen. Specifically, MutGen consists of three main components: (1) extracting and incorporating code summarization into prompts, (2) applying the fixing step to repair invalid test cases, and (3) extracting and incorporating mutation feedback into prompts. In this study, we disable one component at a time, yielding three variants: MutGen-S (using the original code comments instead of summarization), MutGen-F (omitting the fixing step), and MutGen-MF (excluding mutation feedback). For each variant, we perform four iterations of test case generation and measure mutation score to assess the impact of each component on the effectiveness of MutGen.
@@
Implementation. We describe important implementation details of MutGen and our evaluation setup. All subjects used in the evaluation are built using the Maven build system. We follow official guidelines to integrate EvoSuite, JaCoCo, and PITest into the Maven pipeline for automated test generation, code coverage measurement, and mutation testing, respectively. Our approach, MutGen, is implemented in Python and built on top of the Mutahunter framework [48]. Specifically, we reuse components from Mutahunter for prompt handling, LLM interfacing, and response processing. We use Ollama [51] to locally deploy the Llama-3.3 70B model [30] with the default configuration, setting the temperature to 0.0 for more deterministic outputs. As our main baseline, EvoSuite officially and stably supports JUnit 4, while its support for JUnit 5 is currently limited and experimental. Therefore, JUnit 4 remains the recommended choice for reliable test generation. To ensure a fair comparison, we adopt JUnit 4 in all experimental settings.
@@ III-B Results and Analysis
On HumanEval-Java, for both line coverage and branch coverage, MutGen outperforms EvoSuite, EvoSuitemut and Genvanilla. On Leetcode-Java, we see the opposite: EvoSuite achieves the highest coverage scores, surpassing both LLM-based techniques. For both datasets, however, differences are relatively small, with values below 3 percentage points. Although MutGen is primarily designed to improve mutation score, the results also demonstrate that it consistently outperforms Genvanilla in terms of line and branch coverage. This supports previous observations  [24] that improvements in mutation score can also lead to enhanced code coverage.
@@
Regarding the mutation score, PITest generates an average of 11 and 19 mutants per subject on HumanEval-Java and Leetcode-Java, respectively, resulting in a total of 1,144 and 1,900 mutants across the two datasets. MutGen achieves mutation scores of 89.5% and 89.1% on the two datasets. Compared to EvoSuite, this represents large improvements of 28.8% and 51.3%, respectively, corresponding to around 20 and 30 percentage points. Compared to Genvanilla, MutGen also significantly improves the mutation score, though to a lesser extent. Additionally, MutGen also surpasses EvoSuitemut, which specifically aims to maximize the mutation score, to an even larger extent. Indeed, surprisingly, we observe that EvoSuite slightly outperforms EvoSuitemut. This might be due to the latter being developed 10 years ago [26] and receiving no further maintenance. For example, we found that the mutation operators used by EvoSuitemut are limited, resulting in poor performance of the generated test cases when evaluated by PITest. Another tentative explanation is that, to ensure fair comparisons, recall that we set the time budget to the maximum used by MutGen, but a larger budget may be necessary for EvoSuitemut.
@@
Furthermore, we report the A12 effect size [7] on mutation scores in the third-to-last column of Table I. For the comparison between EvoSuite and MutGen, the A12 values are 0.759 and 0.899, indicating that MutGen has a 75.9% and 89.9% probability of achieving higher mutation scores than EvoSuite on HumanEval-Java and Leetcode-Java, respectively. Similarly, compared with Genvanilla, the A12 values are 0.650 and 0.734, indicating that MutGen outperforms Genvanilla with probabilities of 65.0% and 73.4% on the same datasets. Additionally, based on the results of the paired-sample Wilcoxon signed-rank test, we found that MutGen achieves statistically significant improvements ($\alpha$ = 0.05) in mutation scores compared with EvoSuite, EvoSuitemut, and Genvanilla. However, there is no statistically significant difference in coverage, as most subjects already achieve near 100% coverage across all approaches.
@@
In terms of overhead, MutGen consumes fewer tokens on HumanEval-Java but more tokens on Leetcode-Java. The LeetCode-Java subjects consist of solutions to medium- and hard-level algorithmic problems, with more complex logic than those in HumanEval-Java. As a result, PITest generates more mutants for them (19 vs. 11), and Llama-3.3 requires more test cases to kill all mutants. Compared with Genvanilla, EvoSuite, and EvoSuitemut, in Leetcode-Java, MutGen consumes more tokens while achieving higher mutation scores as well as better line and branch coverage. In contrast, on HumanEval-Java, MutGen consumes fewer tokens than EvoSuite and EvoSuitemut while still achieving superior performance in both coverage and mutation score. For the baseline approaches (i.e., EvoSuite and EvoSuitemut), a possible explanation for their higher token consumption on HumanEval-Java is that these programs contain more lines of code on average, whereas LeetCode-Java programs exhibit higher cyclomatic complexity. Longer programs tend to yield larger, more verbose test suites generated by these search-based techniques, resulting in increased token usage at the test level. Interestingly, we observe an opposite trend for our LLM-based approach. On LeetCode-Java, MutGen consumes more tokens than on HumanEval-Java, despite the latter having longer programs on average. We hypothesize that this behavior is related to the use of mutation feedback in the prompt. Specifically, LeetCode-Java has more mutants per subject, resulting in more uncovered or live mutants included in the mutation feedback provided to the LLM. As a consequence, the prompt becomes richer and more informative, encouraging the LLM to generate more extensive test cases to kill these mutants, which, in turn, leads to higher token consumption. However, differences of that scale in the number of tokens do not have practical implications.
@@
Overall, the results suggest that MutGen significantly improves mutation scores relative to alternatives, with large differences, at the expense of slightly lower code coverage than EvoSuite for the most complex subject (Leetcode-Java). This is expected, as EvoSuite is designed to maximize coverage. For the other subject (HumanEval-Java), the reverse trend is observed. Even when EvoSuite is specifically configured as EvoSuitemut to maximize mutation scores [26], it performs slightly worse than EvoSuite on both datasets. Therefore, to optimize mutation scores and, thus, fault detection, we recommend using MutGen, though the impact on coverage is unclear and likely negligible in practice.